The sum of a​ student's three scores is 213. If the first is 21 points more than the​ second, and the sum of the first two is 3 more than twice the​ third, what was the first​ score?
Answer #1


& Let x,y,z are the marks of studente Its sum is 213. afyft=213 to.. than Given, first is more than 21 points second x = 21ty put 2=to in equation o 2+ yf 70 = 213 afy = 213-70 afy = 143 + ☺ From @, x-y = 21 . Solving above equatióne x+y = 143 x =21 2